OpenStudy (gabylovesyou):

@phi Hayden is a manager at a landscaping company. He has two workers to landscape an entire park, Cody and Kaitlyn. Cody can complete the project in 8 hours. Kaitlyn can complete the project in 6 hours. Hayden wants to know how long it will take them to complete the project together. Write an equation and solve for the time it takes Cody and Kaitlyn to complete the project together. Explain each step.

OpenStudy (gabylovesyou):

@phi

OpenStudy (gabylovesyou):

ik its 1/8 + 1/6 = 1/x

OpenStudy (phi):

yes, if x is the unknown time

OpenStudy (gabylovesyou):

yes... now.. how do i solve this ?

OpenStudy (phi):

I would add ⅛ + 1/6 you can use a common denominator of 48 or better 24

OpenStudy (phi):

or a calculator

OpenStudy (gabylovesyou):

ok so i have to find out for both what number mulitplies to that to get 24 ?

OpenStudy (phi):

yes, 3 and 4

OpenStudy (gabylovesyou):

so 3/24 and 4/24

OpenStudy (phi):

yes

OpenStudy (gabylovesyou):

7/24

OpenStudy (phi):

yes, so 7/24= 1/x you are allowed to "flip" both sides

OpenStudy (gabylovesyou):

how do i do this ? ;/

OpenStudy (phi):

the fast way is flip each side for example if you had ½ = 4/8 then 2/1 = 8/4

OpenStudy (phi):

the slow way is multiply both sides by x to get x*7/24 = 1 then multiply both sides by 24/7 to get x= 24/7

OpenStudy (gabylovesyou):

24/7 = x/1

OpenStudy (phi):

yes. and x/1 is just x (divide by 1 does not change anything) that is the answer in hours you could write it as a mixed fraction 3 3/7 hours or as a fraction 3.42857 hours or as 3 hours 26 minutes (rounded to the nearest minute)

OpenStudy (gabylovesyou):

why 3 3/7 ?

OpenStudy (phi):

3 hours and 3/7 of an hour it makes a bit more sense (people don't really understand 24/7 hours)

OpenStudy (phi):

if you divide 7 into 24 you get 3 remainder 3 and you write it as 3 and 3/7 or just 3 3/7